OpenStudy (anonymous):

PLEASE HELP! Find the area of the surface obtained by rotating the curve x = 3 e^{2y} from y=0 to y = 1 about the y-axis.

OpenStudy (amistre64):

its always nice to be able to draw a picture for these as well

OpenStudy (amistre64):

it should pretty much be the same as lnx

OpenStudy (amistre64):



OpenStudy (amistre64):

this is just finding the areas of alot of circles from y=0 to y=1

OpenStudy (amistre64):

do you recall the formula for the area of a circle?

OpenStudy (anonymous):

Pi r^2

OpenStudy (amistre64):

do we want surface area of the rotation? or volume of the rotation? the question is not clear in that

OpenStudy (anonymous):

Surface area I think

OpenStudy (amistre64):

oh, i was doing volume :) same concepts, just a different approach

OpenStudy (amistre64):

\[\int 2\pi x\ ds\] where ds = sqrt(1+x'^2) dy